OpenStudy (anonymous):

I need a tutor. Can someone spend like 1 hour helping me with my hw please.

OpenStudy (anonymous):

i can for a bit

OpenStudy (anonymous):

just depends on what your doing i guess

OpenStudy (anonymous):

Same here if need be.

OpenStudy (anonymous):

Im doing perfect trinomials

OpenStudy (anonymous):

----+42z+49

OpenStudy (anonymous):

As in \[(u + v)^{2} = u^{2} + 2 u v + v^{2} ?\]

OpenStudy (anonymous):

the directions say supply the missing term to make each expression a perfect square trinomial

OpenStudy (anonymous):

Suppose the first term is\[a^{2} z^{2}.\] \[\sqrt{a^{2} z^{2} } = az,\]\[\sqrt{49} = 7,\] 2(az)(7) = 14az = 42 z. Solve for a to get a = 3. The blank-filler is \[(3z)^{2},\]or \[9 z^{2}.\]

OpenStudy (anonymous):

You can FOIL to verify that \[(3z + 7)^{2} = 9z^{2} + 42z + 49.\]

OpenStudy (anonymous):

Thank you. Im getting off whenever youre back if you have time could you please help me?

OpenStudy (anonymous):

I can help you some more.

OpenStudy (anonymous):

Yeah cool thanks. But im going out right now.

OpenStudy (anonymous):

See ya :^)